A sample of propane gas [C3H8(g)] in a sliding-piston sampling device (bomb) has a volume of 1.500 L at a temperature of 212.21 K. What temperature in kelvin must the propane be stored at if the volume is increased to 2.500 L
1
Expert's answer
2020-04-24T13:34:29-0400

PV = RT

P = RT / V = 8.314 × 212.21 / 1.5 = 1176.21 kPa

T = PV / R = 1176.21 × 2.5 / 8.314 = 353.68 K

Answer: 353.68 K.